OpenStudy (anonymous):

use only Boolean algebra to simplify the Boolean expression F. (first determine f1 and f2, then simplify f1 +f2) showing all steps f1 = ywx' + yw'x' f2 = (y' + xw)' F(w, x,y) = F1+F2. please assist

OpenStudy (anonymous):

F1 = ywx' + yw'x' = yx'(w+w') = yx' F2 = (y'+xw)' = (y')'.(xw)' = y(x'+w') = yx' + yw' Therefore : F1 + F2 = yx' + yx' + yw = yx' + yw

OpenStudy (anonymous):

Many tnx

OpenStudy (anonymous):

Anytime.. . :)